A ray of light is incident in air at an angle of 40° to the normal to one face of a 60° glass prism.


   Calculate the angle through which the ray has been deviated by the time it emerges from the prism.


1
Expert's answer
2021-03-09T15:36:15-0500

Angle of prism "A=60^{\circ}"

Angle of incidence "i_1=40^{\circ}"

Let angle of deviation be "\delta"


As we know "r_1=r_2=\dfrac{A}{2}=30^{\circ}"

Using snell's law-

 Refractive index "\mu=\dfrac{sin i_1}{sin r_1}=\dfrac{sin 40}{sin30}=1.2855"



Angle of deviation is given by-

"\delta=(\mu-1)A"

"=(1.2855-1)\times 60\\=0.2855\times 60=17.13^{\circ}"